The equation of a line is given by y=(4/50)x-12
Find the equation of a line perpendicular to this line that passes through the point (8,15).
Find the equation of a line parallel to this line that passes through the point (1,1).
Find the equation of a line perpendicular to this line that passes through the point (6,3).

Answer by richwmiller(17219)

I'll give you some hints.
y=(4/50)x-12
says the slope is 4/50=2/25
an equation parallel has the same slope.
an equation perpendicular has a slope of -25/2
parallel y=2/25x+b
perpendicular y=-25/2x+b
plug in your x's and y's to find b
then rewrite the equation with that b
y=-25/2x+b
(8,15)
15=-25/2(8)+b
y=-25/2+b
(6,3)
3=-(25/20*6+b
y=2/25x+b
(1,1)
1=2/25(1)+b
then rewrite the equation with that b